OpenStudy (loser66):

Find horizontal asymptote of \(f(x) =5x^3e^{x}\) Please, help

OpenStudy (loser66):

@SolomonZelman

OpenStudy (loser66):

I would like to know 1) the name of the property: \(lim_{x\rightarrow-\infty} f(x) =\lim_{x\rightarrow \infty}f(-x)\) 2) Is there any way else to find the horizontal asymptote without using that property??

OpenStudy (solomonzelman):

By definition, \(\color{black}{f(x)=\beta }\) is horizontal asymptote iff \(\color{black}{x\to \pm \infty}\) at \(\color{black}{f(x)=\beta}\).

OpenStudy (solomonzelman):

So, going from the definition, your horizontal asymptote is \(\color{black}{f(x)=0}\), because \(\color{black}{x\to -\infty }\) at that point. (Although, this is sure not like the "typical" asymptote case.)

OpenStudy (loser66):

Logic, please.

OpenStudy (loser66):

How can you get 0?

OpenStudy (solomonzelman):

\(\color{black}{\lim_{x\to-\infty}5x^3e^x=0}\)

OpenStudy (loser66):

We have \(lim_{x\rightarrow \infty}\dfrac{x^n}{e^x}=0\) So, we have to apply the property 1) above to get that form. Right?

OpenStudy (tackopp):

Hey, how do you ask questions? Really need help with a question!

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \lim_{x\to-\infty}5x^3e^x=\lim_{x\to\infty}\frac{-5x^3}{e^x}=-5\lim_{x\to\infty}\frac{x^3}{e^x}}=0\) if you want it in that form, ... but even without this, we can surely tell the limit is 0.

OpenStudy (loser66):

Common, I need know the name of that property. Not just "trust me!!, we can surely get it"

OpenStudy (solomonzelman):

That the line \(\color{black}{\displaystyle y=b}\) is a horizontal asymptote of \(\color{black}{\displaystyle f}\) if \(\color{black}{\displaystyle \lim_{x\to \pm \infty}f(x)=b}\).

OpenStudy (solomonzelman):

You need the name of this property? I don't have it ...

OpenStudy (loser66):

Sometimes, you have y = b is a horizontal asymptote of f from both sides \(\pm \infty\) Sometimes, we have just one way. not both

OpenStudy (solomonzelman):

Yes, that is true ...

OpenStudy (loser66):

And there is no way to find the limit when x approaches to -infinitive, we have to turn it to +infinitive and change the f(x) to f(-x) Unfortunately, not all of profs give the students that property.

OpenStudy (solomonzelman):

oh you want the "official" way of solving that limit?

OpenStudy (loser66):

Yes, please

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \lim_{x\to -\infty} 5x^3e^x}\) \(\tiny \\[0.9em]\) Substitution: \(\color{black}{\displaystyle x=-p}\)\(\tiny \\[0.2em]\) Then, \(\color{black}{\displaystyle 5x^3=-5p^3{\tiny~}}\) and \(\color{black}{\displaystyle e^x=e^{-p}=\frac{1}{e^p}{\tiny~}}\) (by direct substitution).\(\tiny \\[0.9em]\) Also, as \(\color{black}{\displaystyle x\to -\infty }\), then \(\color{black}{\displaystyle -p\to -\infty }\), and therefore \(\color{black}{\displaystyle p\to +\infty }\). Thus, your new limit is \(\color{black}{\displaystyle \lim_{p\to \infty} \frac{-5p^3}{e^p}=-5\lim_{p\to \infty} \frac{p^3}{e^p}}\) and knowing that \(\color{black}{\displaystyle \forall n,m\in\mathbb{R}\left[m\times \lim_{p\to \infty} \frac{p^n}{e^p}=0\right]}\), we get \(\color{black}{\displaystyle -5\lim_{p\to \infty} \frac{p^3}{e^p}=0}\).

OpenStudy (solomonzelman):

((The most important thing in this "official" is the substitution, really, and that we apply the property of the limit at infinity, of a polynomial over exponentially increasing function ... any more rigorousness is just a spice))

OpenStudy (loser66):

Thank you so much. I got it.

OpenStudy (solomonzelman):

Just didn't know what exactly you were asking to back up:) YW

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \lim_{z\to \infty} \frac{e^x}{\Gamma(z)}=0}\) :)

OpenStudy (solomonzelman):

jk

OpenStudy (loser66):

oh no, gamma function!!

OpenStudy (solomonzelman):

You don't like the gamma function?